7. Show that the equation f(x) = x^3 + 3x^2 - 9x + 7 = 0 has a solution for some x is E(-6; -5).
Apply Newton’s method with an initial guess x0 = -5 to find x2.


8. Find the intervals of increase and decrease of the function x2e^-2x.


9. Sketch the graph of the curve y = x3 + 3x2 - 9x + 7. Be sure to find the intervals of increase,
decrease and constant concavity and all local extremes and inflection points and all intercepts.
If you don’t know the intercept(s) exactly, use Newton’s method to get it accurarte to one
tenth.
Answer #1

Solu, Giren us. equation fees x + 3x² - gx +7=0 Newtons Method have In a fren) Xntia f(on) 2881.2 differenkieling eq en watBy newhon methods U2 X, f(xi) 5(4) 1-26.24 CE X = -3.3125 4.035 2 = -3.3125 +6.50 X = 3.1875 of. -0) -27, - 2 - 2x - 2x 2x -8 1 to Intervals signs decrease tit (-0,0) (0,1) (1,0) Increase Intrete decrease ti 11) Ctit Increase 111) g Given us. x + 3xand for extreme values. Again differentiating eq (1) wird a alors p(m)6x +6 me # At x = -3 le - 34 1(-3) = -18+6 =-12 (nega